# Novel SiC UV Instrumentation Development with Potential Applications for the Habitable Worlds Observatory

### v1

@@ 5 Current Testing of Photodetector Properties and Near Term Future Work
Figure 6: Characterization of large format SiC photodiodes for ECI effort including (a) broad spectrum quantum efficiency characterization and (b) dark current measurement. Additional development work is expected to improve QE values though current values already satisfy Pandora-SiC goals, while dark current is an upper bound on the true value given measurement limitations.
Similarly, initial testing of the I-V curves and QE has been conducted for photodiodes designed under an SBIR program that can be thought of as a precursor to the PReSSiC photodiodes. Under this program, three different diodes that had been fabricated were tested. The first is an $n^{+}\text{-}p$ diode, the second an $n^{-}\text{-}p$ diode and the third, a Schottky diode. The $n^{+}\text{-}p$ and $n^{-}\text{-}p$ diodes are both 40-$\mu$m pitch diodes with exposed active areas of about $172.5\times 10^{-6}\,\text{mm}^{2}$. The Schottky diode is also a 40-$\mu$m pitch diode with an exposed active area of $270\times 10^{-6}\,\text{mm}^{2}$. I-V curves were constructed by taking current and voltage values under illumination from 200-350nm, in 5nm increments. Quantum efficiency values were also calculated for all three diodes in the same wavelength range.
QE values for the $n^{+}\text{-}p$ diode peaked at 0.8 at 260nm, were above 0.6 from 240-280nm and were above 0.2 from 200 to 330nm. QE values for the $n^{-}\text{-}p$ diode peaked at a little bit above 0.8 at 260nm, were nearly at or above 0.6 from 230 to 290nm and were above 0.2 for a similar wavelength range as the $n^{+}\text{-}p$ diode. QE values for the Schottky diode peak from 250-260nm at nearly 0.7, were 0.6 or higher from 230-270nm and were above 0.2 from 200-320nm. Optical I-V curves exhibited a similar shape and general trend across all three diodes, though we note there was an offset in voltage where current approached zero which was likely due to an instrumental effect. Using the dynamic resistance calculated form the I-V curves and the QE values, we were able to calculated Specific Detectivity (D) and Noise Equivalent Power (NEP) values for each of the diodes over the 200-350nm wavelength range.
@@
The values for NEP and D for the diodes are given in Figure 7(a) and 7(b). The three diodes exhibit broadly the same relationships and trends versus wavelength, with the $n^{+}\text{-}p$ and $n^{-}\text{-}p$ diodes yielding slightly lower NEP values. The 230-300 nm range possess the most favorable noise characteristics across all three diodes with a drop off in sensitivity towards longer wavelengths. While these measurements are important initial characterization steps for development of these detectors, we note they are minimum bounds on sensitivity of the diodes (put another way, they are an overestimate of noise). This is because measurements of the I-V curve were done at coarse spectral resolution, and more highly sampled measurements where current drops to zero are likely to yield even lower dynamic resistance values for the diodes. In addition to the measurement improvements in both ECI and the SBIR/PReSSiC characterization efforts, we note that in both cases these are early stage diodes in the development process. We expect that as design and fabrication processes mature, NEP and D values are likely to improve.
Future characterization and testing includes both improvements to both testing methodology and hardware as well as additional tests that correspond to future development needs. Enabling very low current measurements and carrying these measurements out at highly sampled voltage values near a current minima are high near-term priorities for our team. In the future, as additional diodes are available for testing in our upcoming production runs, we will conduct additional tests to optimize diode parameters prior to focal plane array/spectrometer design. Once that is achieved, and a focal plane array/spectrometer is fabricated, our team will utilize pre-existing optical testbeds for the Pandora-SiC and PReSSiC projects to carry out instrument testing.
@@ 6 A promising detector material for the Habitable Worlds Observatory
The Habitable Worlds Observatory (HWO) is recommended in the most recent astrophysics Decadal Survey [1] to support broad coverage of 100 nm - 2.5 $\mu$m wavelength through a variety of instrument modes, including high-contrast direct imaging of circumstellar material and bodies through coronagraphy, high angular resolution imaging, integral field spectroscopy, multi-object spectroscopy, and single-object spectroscopy. NUV capability of each of these modes is required to drive advances in our understanding of exoplanet atmospheres, stars, interstellar medium, galaxies, and solar system science. At this early stage in the mission concept design, the HWO Technology Maturation Project Office has adopted instruments from the LUVOIR-B mission concept [27] as a starting design and is additionally considering a UV coronagraph instrument (G. Arney, private communication). There are no requirements yet for these instruments, but based on the LUVOIR report, there is a clear need for large format arrays that are buttable on three sides and have high quantum efficiency, low dark current, and low read noise. Some instrument modes may not require large format arrays but may have more stringent requirements on efficiency and noise. The LUVOIR telescope was planned to be actively controlled to an operating temperature of 270 K and had set limits on mass and power for different observatory segments [28]. SiC detectors have potential to satisfy or achieve many of these requirements and desired properties.
@@
The key scientific motivation for the UV coronagraph is to search for ozone on potentially habitable exoplanets. Ozone is a significant biosignature on the Earth, because it is generated from the photolysis of oxygen by UV sunlight. The Hartley-Huggins band of ozone between 200-350 nm is much stronger than the Chappuis band near 600 nm and would be a sensitive tracer of ozone at levels significantly lower than present day values. This is key because oxygen and consequently ozone are likely to have varied significantly over Earth’s history, and over much of the Earth’s past are likely to have been at much lower levels than present day values. For example, prior to the great oxygenation event, around 1-2 billion years ago, the abundance of oxygen in Earth’s atmosphere is likely to have been only 0.1% of present-day levels [29]. Current oxygen and ozone levels are a relatively recent feature of Earth’s atmosphere, and consequently, there is increasing recognition that sensitivity to smaller abundances of the two molecules is desirable when probing whether a world may be habitable. Thus, NUV spectroscopy is a powerful probe of smaller amounts of ozone, that may enable characterization of a planet’s habitability even when key biosignature gases are at abundances significantly lower than present day Earth’s abundances.
@@
To demonstrate the promise of SiC detectors for HWO, we describe a use case of measuring ozone in an Earth-like exoplanet with the UV coronagraph in comparison to other potential detector material. We assume the UV coronagraph would be operable over the 225-450 nm range, but with a smaller instantaneous bandpass and optimized for 250 nm. We focus on the 210-350 nm range that captures contributions from the ozone signatures in the longer wavelength portion of that region and that aligns with wavelength regions of focus for the two projects detailed in previous sections. We assume we are observing a potential exo-Earth candidate at a distance from Earth that corresponds to the Tau Ceti system. Multiple HWO target lists indicate that such a system would have favorable observational properties for HWO[30], including completeness of systems’ habitable zone[31] (though we note there are likely even more favorable systems that are closer to Earth).
@@
We take atmospheric profiles for the simulated planet from Whole Atmosphere Community Climate Model—WACCM6 runs that simulate atmospheres relevant to previously inferred states of the Earth atmosphere prior to the present day[32]. We specifically choose oxygen and corresponding ozone profiles from the 1 and 10% O2 PAL simulations - these scenarios match the range of expected abundances for much of the Proterozoic and Phanerozoic eras, which cover more than half of Earth’s history. We note that while corresponding values are taken for ozone and oxygen, we do not self-consistently model the chemical abundances of other low or trace abundance atmospheric constituents. Most of these gases are not abundant enough to impact spectra, particularly in the near-ultraviolet, and are consequently not as critical to the goals of our simulation for this paper. However, we do intend to incorporate chemically self-consistent atmospheres in future work. Spectra of the planet is modeled using the Planetary Spectrum Generator (PSG) [33]. Using PSG, we simulate spectra of the system as viewed edge on with the planet at quadrature, with sampling techniques and radiative transfer details given in previous work [34].
(a) Simulated HWO Spectra for a 1% O2 PAL Case
(b) Simulated HWO Spectra for a 10% O2 PAL Case
@@
Figure 8: Simulated HWO spectra for an exo-Earth with atmospheric profiles corresponding to 1 and 10% O2 PAL cases. NEP values are taken for SiC versus other material and demonstrate the unique ability of SiC to enable access to the NUV signature on 3 day timescales.
@@
Target SiC instrument parameters are taken from previous and current work and incorporated in PSG’s instrument module. An NEP value of 1E19 for SiC is taken from earlier SiC work[3] and is scaled based upon ${}^{*}/$NEP values in figure 7, while an NEP value of 1E17 is used as a comparison value for other detectors such as AlGaN and Si detectors[3]. The coronagraph model is based upon the PSG’s LUVOIR Extreme Coronagraph for Living Planetary Systems (ECLIPS) coronagraph model, with a modified 6.7m inscribed diameter for the telescope, a spectral resolution of R=7 and an exo-zodi level of 3 assumed for the system. A QE value of 0.8, emissivity of 0.2 and telescope temperature of 270K are taken from requirements for LUVOIR, and are all targets for current SiC development. Figure 8(a) and 8(b) show results for the two simulated cases (1 and 10% O2 PAL) with a comparison of the performance of the different detectors. Total observation time for both cases is 3 days, which is chosen to achieve SNR of at least 3 for the increased flux due to ozone at the longer wavelengths in the better performing detector, which in both cases in SiC. The integration time and detector properties will vary as requirements are clarified, simulation parameters are optimized and fabrication and detector development progress is made. However, the general relationship from the two simulations is clear and robust - the potential dark current advantages that SiC may offer from similar packaging configurations is likely to yield more robust detections of potential signatures in the NUV, by at least an order of magnitude in SNR. Of note is that there are potential degeneracies with rayleigh scattering and aerosols in this spectral region, but there should be differences in the spectral signatures between those confounding sources versus ozone that the improved SNR may be able to leverage.
@@
The improved noise performance that SiC may offer would translate to other potential UV instruments that HWO may baseline. While the format and design of such detectors would vary based upon the specific use cases, our team is examining potential applicability for different instruments. The improved noise performance and other characteristics of future SiC detectors would also be useful for other observational platforms in astrophysics and other disciplines. All of these potential use cases will benefit from additional technology and engineering advancement work that is informed by science modeling, including the work that our team is carrying out and will carry out in the near future as part of our ongoing projects.
@@
The authors declare that there are no financial interests, commercial affiliations, or other potential conflicts of interest that could have influenced the objectivity of this research or the writing of this paper.

### v2

@@ 5 Current Testing of Photodetector Properties and Near-Term Future Work
Figure 6: Characterization of large format SiC photodiodes for Pandora+SiC effort including (a) broad spectrum quantum efficiency characterization and (b) dark current measurement. Additional development work is expected to improve QE values though current values already satisfy Pandora+SiC goals, while dark current is an upper bound on the true value given measurement limitations.
Similarly, initial testing of the I-V curves and QE has been conducted for three different photodiodes designed under an SBIR program that are is precursor to the PReSSiC program. The first is an $n^{+}\text{-}p$ diode, the second an $n^{-}\text{-}p$ diode and the third, a Schottky diode. The $n^{+}\text{-}p$ and $n^{-}\text{-}p$ diodes are both 40-$\mu$m pitch diodes with exposed active areas of about $172.5\times 10^{-6}\,\text{mm}^{2}$. The Schottky diode is also a 40-$\mu$m pitch diode with an exposed active area of $270\times 10^{-6}\,\text{mm}^{2}$. I-V curves were constructed by measuring current values, using trans resistance amplifiers without integration, under illumination from 200-350nm, in 5nm increments, at varying bias values. Quantum efficiency values were also calculated for all three diodes in the same wavelength range.
QE values for the $n^{+}\text{-}p$ diode peaked at 0.8 at 260nm, were above 0.6 from 240-280nm and were above 0.2 from 200 to 330nm. QE values for the $n^{-}\text{-}p$ diode peaked at a little bit above 0.8 at 260nm, were nearly at or above 0.6 from 230 to 290nm and were above 0.2 for a similar wavelength range as the $n^{+}\text{-}p$ diode. QE values for the Schottky diode peak from 250-260nm at nearly 0.7, were 0.6 or higher from 230-270nm and were above 0.2 from 200-320nm. Optical I-V curves exhibited a similar shape and general trend across all three diodes, though we note there was an offset in voltage where current approached zero which was likely due to an instrumental effect. Using the dynamic resistance calculated from the I-V curves and the QE values, we were able to calculate Specific Detectivity (D) and Noise Equivalent Power (NEP) values for each of the diodes over the 200-350 nm wavelength range.
@@
The values for NEP and D for the diodes are given in Figure 7(a) and 7(b). The three diodes exhibit broadly the same relationships and trends versus wavelength, with the $n^{+}\text{-}p$ and $n^{-}\text{-}p$ diodes yielding slightly lower NEP values. The 230-300 nm range possess the most favorable noise characteristics across all three diodes with a drop off in sensitivity towards longer wavelengths. While these measurements are important initial characterization steps for development of these detectors, we note they are minimum bounds on sensitivity of the diodes. This is because measurements of the I-V curve were done at coarse spectral resolution, and more highly sampled measurements where current drops to zero are likely to yield even lower dynamic resistance values for the diodes. In addition to the measurement improvements in both Pandore+SiC and the SBIR/PReSSiC characterization efforts, we note that in both cases these are early-stage diodes in the development process. We expect that as design and fabrication processes mature, NEP and D values will improve.
Future characterization and testing include improvements to both testing methodology and hardware as well as additional tests that correspond to future development needs. Enabling very low current measurements, including signal integration for correlated double sampling, and carrying measurements out at highly sampled voltage values near a current minima is a high near-term priority. In the future, as additional diodes from upcoming production runs are available for testing, we will conduct additional tests to optimize diode parameters prior to focal plane array/spectrometer design. Once that is achieved, and a focal plane array and spectrometer is fabricated, our team will utilize pre-existing optical testbeds for the Pandora+SiC and PReSSiC projects to carry out instrument testing.
@@ 6 A promising detector material for the Habitable Worlds Observatory
The Habitable Worlds Observatory (HWO) was recommended by the most recent astrophysics Decadal Survey [1] to support broad coverage of 100 nm - 2.5 $\mu$m wavelength through a variety of instrument modes, including high-contrast direct imaging of circumstellar material and bodies through coronagraphy, high angular resolution imaging, integral field spectroscopy, multi-object spectroscopy, and single-object spectroscopy. NUV capability of each of these modes is required to advance our understanding of exoplanet atmospheres, stars, interstellar medium, galaxies, and solar system science. At this early stage in the mission concept design, the HWO Technology Maturation Project Office has adopted instruments from the LUVOIR-B mission concept [35] as a starting design and is additionally considering a UV coronagraph instrument (G. Arney, private communication). There are no requirements yet for these instruments, but based on the LUVOIR report, there is a clear need for large format arrays that are buttable on three sides and have high quantum efficiency, low dark current, and low read noise. Some instrument modes may not require large format arrays but may have more stringent requirements on quantum efficiency and noise. The LUVOIR telescope was planned to be actively controlled to an operating temperature of 270 K and had set limits on mass and power for different observatory segments [36]. SiC detectors have the potential to satisfy or achieve many of these requirements and desired properties. In the rest of this section we focus on a UV coronagraphy application for SiC.
@@
The key scientific motivation for the UV coronagraph is to search for ozone on potentially habitable exoplanets. Ozone is a significant biosignature on the Earth, because it is generated from the photolysis of oxygen by UV sunlight. The Hartley-Huggins band of ozone between 200-350 nm is much stronger than the Chappuis band near 600 nm and would be a sensitive tracer of ozone at levels significantly lower than present day values. This is key because oxygen and consequently ozone are likely to have varied significantly over Earth’s history, and over much of the Earth’s past are likely to have been at much lower levels than present day values. For example, prior to the great oxygenation event around 1-2 billion years ago, the abundance of oxygen in Earth’s atmosphere is likely to have been only 0.1% of present-day levels [37]. Current oxygen and ozone levels are a relatively recent feature of Earth’s atmosphere, and consequently, there is increasing recognition that sensitivity to smaller abundances of the two molecules is desirable when probing whether a world may be habitable. Thus, NUV spectroscopy is a powerful probe of smaller amounts of ozone, that may enable characterization of a planet’s habitability even when key biosignature gases are at abundances significantly lower than present day Earth’s abundances.
@@
To demonstrate the promise of SiC detectors for HWO, we describe a use case of measuring ozone in an Earth-like exoplanet with the UV coronagraph in comparison to other potential detector material like Si and AlGaN. We assume the UV coronagraph would be operable over the 225-450 nm range, but with a smaller instantaneous bandpass and optimized for 250 nm. We conduct simulations over the 210-350 nm range in order to capture the Hartley-Huggins ozone band and align with wavelength regions of focus for the two SiC development projects detailed in previous sections. Future SiC development projects could optimize for different wavelength regions. We assume we are observing a potential exo-Earth candidate at a distance from Earth that corresponds to the Tau Ceti system (3.7 pc). Multiple HWO target lists indicate that such a system would have favorable observational properties for HWO[38], including accessibility of the systems’ entire habitable zone[39] (though we note there are likely even more favorable systems that are closer to Earth).
@@
We assume atmospheric profiles for the simulated planet from Whole Atmosphere Community Climate Model—WACCM6 runs that simulate atmospheres relevant to Early Earth[40]. We specifically choose oxygen and corresponding ozone profiles from the 1 and 10% O2 PAL simulations - these scenarios match the range of expected abundances for much of the Proterozoic and Phanerozoic eras, which cover more than half of Earth’s history. We note that while corresponding values are taken for ozone and oxygen, we do not self-consistently model the chemical abundances of other low or trace abundance atmospheric constituents. Most of these gases are not abundant enough to impact spectra in the simulated 210-350 nm range. However, we do intend to incorporate chemically self-consistent atmospheres in future work. The planet’s spectrum is modeled using the Planetary Spectrum Generator (PSG) [41]. Using PSG, we simulate spectra of the system as viewed edge on with the planet at quadrature, with sampling techniques and radiative transfer details given in previous work [42].
@@
Figure 8: Simulated HWO spectra for an exo-Earth with atmospheric profiles corresponding to 1% and 10% O2 PAL cases. We assumed an integration time of three days and a resolving power R=7 (within the spectral range). NEP values are taken for SiC (black squares and error bars) versus other material (orange error bars). The underlying expected contrast ratio across the Hartley-Huggins ozone band is shown in gray for reference. This figure demonstrates the unique ability of SiC to enable access to the NUV ozone signature on 3-day timescales.
@@
Target SiC instrument parameters are taken from previous and current work and incorporated in PSG’s instrument module. An NEP value of 1019 W/$\sqrt{Hz}$ for SiC is taken from earlier SiC work[3] and is scaled based upon ${}^{*}/$NEP values in figure 7, while an NEP value of 1017 W/$\sqrt{Hz}$ is used as a comparison value for other detectors such as AlGaN and Si detectors[3]. The coronagraph model is based upon the PSG’s LUVOIR Extreme Coronagraph for Living Planetary Systems (ECLIPS) coronagraph model, with a modified 6.7m inscribed diameter for the telescope, a spectral resolution of R=7 and an exo-zodi level of 3 assumed for the system. A QE value of 0.8, emissivity of 0.2 and telescope temperature of 270K are taken from requirements for LUVOIR, and are all targets for current SiC development. Figure 8 shows results for the two simulated cases (1 and 10% O2 PAL) with a comparison of the performance of the different detectors. Total observation time for both cases is 3 days, which is chosen to achieve SNR of at least 3 for the increased flux due to ozone at the longer wavelengths in the better performing detector, which in both cases in SiC. The integration time and detector properties will vary as requirements are clarified, simulation parameters are optimized and fabrication and detector development progress is made. However, the general relationship from the two simulations is clear and robust - the potential dark current advantages that SiC may offer from similar packaging configurations is likely to yield more robust detections of potential signatures in the NUV, by at least an order of magnitude in SNR. There are potential degeneracies between ozone and Rayleigh scattering and aerosols in this spectral region that the improved SNR offered by a superior detector could help resolve. This simulation also does not capture the advantage of visible blindness during measurement which would lead to further improved SNR and higher reliability in measurement.
@@
The improved noise performance that SiC may offer would translate to other potential HWO instruments as well as other observational platforms in astrophysics and other disciplines. While the format and design of such detectors would vary based upon the specific use cases, our team is examining potential applicability for a variety of science cases to inform additional technology and engineering advancement work of SiC detectors.
@@
The authors declare that there are no financial interests, commercial affiliations, or other potential conflicts of interest that could have influenced the objectivity of this research or the writing of this paper. While the sensor and detector technologies and scientific work described in this article will be presented in proposals for future competitive calls and business opportunities, this has not influenced the dissemination of the results of the past and ongoing work in this article.
Code, Data, and Materials Availability
Data for this study is displayed in figures in the body of the paper. Underlying config files and additional data for the spectra in the HWO simulation section are available upon request. Source I-V and QE data related to NEP and D are also available upon request.